# F-95000 Cergy-Pontoise, France

### v1

@@ 5.3 Analytic Queries and their Consistent Answers
In the literature Ullman , an analytic query is a query involving an aggregate function among $count$, $min$, $max$, $sum$, and possibly a group-by clause. Formally, given a data warehouse over a star schema, an analytic query $\mathcal{AQ}$ has one of the following two forms:
@@
where $\varphi$ is the join of the fact table $F$ with all dimension tables $D_{1},\ldots,D_{n}$, $X$ is a relation schema, and for $j=1,\ldots,p,$ $aggr$ is an aggregate function such as $count$, $min$, $max$, $sum$ and $M_{i}$ is a measure attribute.
@@
  • If $\mathcal{AQ}$ involves no group-by clause, ${\sf Ans}(\mathcal{AQ})$ is the value of the aggregation evaluated over the set of all tuples in $\varphi$ satisfying the condition $\Gamma$. We notice that the aggregate may involve no attribute, when expressed as $count(*)$.
@@
Now, if the underlying data warehouse does not satisfy all functional dependencies, this traditional semantics of answer to analytic queries has to be revisited. When no dependency of the form $\mathbf{K}\to M_{i}$ is considered, this has been done in ArenasBCHRS03 , in the case of queries with no group-by clause and in FuxmanFM05 when a group-by clause occurs in the query. As shown in DixitK19 , in either case, the ‘repair semantics’ of the consistent answer to an analytic query consists intuitively in producing an interval in which aggregate values fall, when answering the query in any repair. The reader is referred to the forthcoming Section 7 for a more precise relationship between these approaches and our work.
In our approach, we follow the same line regarding analytic queries and their consistent answers, and we emphasize that, as in the previous section, all selection conditions are assumed to satisfy the restrictions that are independent and that they involve no attribute from $\mathbf{K}$.
@@
  • If there exists $R$ in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no tuple over $sch(Q)$ satisfying $\Gamma$, then ${\sf C\_ans}(\mathcal{AQ})={\tt NULL}$ if $aggr\neq count$, and ${\sf C\_ans}(\mathcal{AQ})=0$ if $aggr$ is $count$.
@@
  • for every $R$ in ${\sf Rep}(T)$ there exists $d$ in $[glb,lub]$ such that ${\sf Ans}(\mathcal{AQ}^{[R]})$ contains $(x\,d)$,
  • there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})$ and ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})$ respectively contain $(x\,glb)$ and $(x\,lub)$. $\Box$
@@
  • In case (A), a NULL answer means that the consistent answer to the query $Q:~{}{\tt select}$ $\mathbf{K}$ from $T~{}{\tt where}~{}\Gamma$ is empty. In this case the aggregate cannot be computed if different than $count$, and if the aggregate is $count$, then the expected answer is 00. This explains why we return NULL in the former case.
@@
Additionally to consistent answers as defined above, we argue that our approach allows for another kind of consistent answers to analytic queries, relying on the tuples in ${\sf Cons}(\mathcal{T})$, or by Theorem 4.1 on tuples true in every repair. To define this kind of consistent answer, a given analytic query $\mathcal{AQ}:~{}$select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] is associated with a query involving no aggregate, defined by
$Q_{\mathcal{AQ}}:$ select $\mathbf{K},$[$X$]$,M_{i}$ from $T$ where $\Gamma$.
Then, relying on the consistent answer to $Q_{\mathcal{AQ}}$, we define the strongly consistent answer to $\mathcal{AQ}$, denoted by ${\sf C\_ans}^{*}(\mathcal{AQ})$, as the answer to the following analytic query:
select [$X$]$,aggr(M_{i})$ from ${\sf C\_ans}(Q_{\mathcal{AQ}})$ [group-by $X$].
We emphasize that, based on Corollary 1, since the select clause of $Q_{\mathcal{AQ}}$ contains $\mathbf{K}$, ${\sf C\_ans}(Q_{\mathcal{AQ}})$ can be effectively computed, assuming that $m\_Chase(T)$ is available. The important issue of efficiently computing answers to analytic queries is addressed in the next section.
@@
In the context of Example 3, $\mathcal{AQ}_{1}:{\tt select}$ $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$) is an example of analytic query with no group-by clause and involving the aggregate $\max$. The expected result is the interval containing all possible sums of $M_{1}$-values among all tuples defined over $K_{1}K_{2}$ and associated with $A_{1}^{1}$-value $a_{1}$ in all repairs. Referring to Figure 3, we have:
@@
Assuming that $m_{1}$, $m^{\prime}_{1}$ and $m^{\prime\prime}_{1}$ are positive numbers and that $m^{\prime}_{1}\leq m^{\prime\prime}_{1}$, we obtain that ${\sf C\_ans}(\mathcal{AQ}_{1})=[glb,lub]$ and where $glb=m_{1}$ and $lub=2.m_{1}+m^{\prime\prime}_{1}$.
On the other hand, we have $Q_{\mathcal{AQ}_{1}}:~{}$select $\mathbf{K},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$), and it can be seen from Figure 3, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{1}})=\{k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$. Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{1})=m_{1}$.
To illustrate that if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$ and ${\sf C\_ans}^{*}(\mathcal{AQ})=m$, then it does not always hold that $glb\leq m\leq lub$, consider the aggregate $count(*)$ instead of $sum(M_{1})$. Indeed, in this case, ${\sf C\_ans}(\mathcal{AQ})=[3,3]$ and ${\sf C\_ans}^{*}(\mathcal{AQ})=1$.
@@
On the other hand, we have $Q_{\mathcal{AQ}_{2}}:~{}$select $K,A_{1}^{2},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$). It can be seen from Figure 3, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{2}})=\{k_{1}k_{2}b_{1}m_{1}\}$. Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{2})=\{(b_{1},m_{1})\}$.
This example illustrates that if $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$ and $x\,m$ is in ${\sf C\_ans}^{*}(\mathcal{AQ})$, then it does not always hold that $glb\leq m\leq lub$. Indeed, for $m_{1}=1$, $m^{\prime}_{1}=2$ and $m^{\prime\prime}_{1}=3$, we have $glb=3$, $lub=4$ showing that for $x=b_{1}$, $m\not\in[3,4]$. $\Box$
@@ 6 Computing Consistent Answers to Analytic Queries
In this section, we give algorithms for computing consistent answers to analytic queries, first when no group-by clause is present, then when a group-by clause is present. We also consider in this section the case of analytic queries involving a group-by-having clause, whereas the introduction of the clause distinct is the subject of the last sub-section. We emphasize again that the selection conditions in queries are independent and involve no key attributes.
@@ 6.1 Basic Result
Before presenting our algorithms, we state important properties of repairs and analytic queries. To this end, we introduce the following additional notation. Given a star-table $T$ and an analytic query $\mathcal{AQ}:~{}$select [$X$],$\,aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$], let:
$\bullet$ $\Sigma(\mathcal{AQ})=\{\sigma\in m\_Chase(T)~{}|~{}(\mathbf{K}\cup M_{i}\cup sch(\Gamma)\subseteq sch(\sigma))\wedge(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset)\}$
@@
Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}:~{}$select [$X$],$\,aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] an analytic query where $\Gamma$ is independent, i.e., $\Gamma=\Gamma(A_{1})\wedge\ldots\wedge\Gamma(A_{k})$ where $sch(\Gamma)=A_{1}\ldots A_{k}$. If $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})\neq\emptyset$, there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that
@@
    for every $\sigma$ in $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$, there exists $t$ in ${\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{1})$ such that $t.sch(\Gamma)\not\in Sat(\Gamma)$,
@@
    for every $\sigma$ in $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$, there exists $t$ in ${\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{2})$ such that $t.sch(\Gamma)\in Sat(\Gamma)$.
Proof. The proof relies on the fact that for every repair $R$ of $T$ and every $\sigma$ in $\Sigma(\mathcal{AQ})$, ${\sf True}(\mathcal{R})$ contains exactly one tuple $t$ such that $t\in{\sf tuples}(\sigma)$ and $t.K\in\sigma(\mathbf{K})$ (the existence is a consequence of Proposition 4, unicity follows from Proposition 2(1) because $|\sigma(\mathbf{K})|=1$, $\mathbf{K}$ is a key of $R$ and $R\models FD$). As a consequence of Proposition 6, the existence of repairs $R_{1}$ and $R_{2}$ is explicitly shown using the process (P).
Step P1. For every $A_{p}^{k}$ in $sch^{*}(D_{p})\cap sch(\sigma)$ such that $R_{1}$ contains no tuple $k_{p}a_{p}^{k}$ from ${\sf tuples}(\sigma(K_{p}A_{p}^{k}))$
$(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\setminus Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$. If ${\sf tuples}(\sigma(A_{p}^{k}))\subseteq Sat(\Gamma(A_{k}^{p}))$, choose any $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$. It is important to notice that in this case, there always exists at least one attribute $A_{p}^{k}$ such that ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, because otherwise $\sigma$ would be in $\Sigma^{+}(\mathcal{AQ})$.
Step P2. If $M_{i}\not\in sch(\Gamma)$ or if $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma^{+}(\mathcal{AQ})$, choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k_{p})=m$. Otherwise (i.e., $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$), choose an $M_{i}$-value $m$ in $\sigma(M_{i})\setminus Sat(\Gamma(M_{i}))$ for $\varphi_{i}(K_{p})$.
Once all m-tuples in $\Sigma(\mathcal{AQ})$ have been considered, the two steps above are completed by considering all non-processed $K_{i}$- or $\mathbf{K}$-values as done in the generic description of process (P). Consequently, the corresponding set $\varphi$ in process (P) is properly defined.
Step P3. Build up the tuples $t_{\varphi}(\sigma)$ for every $\sigma$ in $m\_Chase(T)$. We notice that
$(a)$ if $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\in Sat(\Gamma)$,
$(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\not\in Sat(\Gamma)$.
2. Regarding the item 2, the repair $R_{2}$ is obtained as above, by changing $(b)$ in Step P1 and Step P2 as follows:
Step P1. $(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\cap Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$. It is important to notice that in this case, every attribute $A_{p}^{k}$ in $sch(\Gamma)$ is such that ${\sf tuples}(\sigma(A_{p}^{k}))\cap Sat(\Gamma(A_{k}^{p}))\neq\emptyset$, because otherwise ${\sf tuples}(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ would not be possible.
Step P2. Choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k)=m$ where $k=\sigma(\mathbf{K})$.
Then, as above, these two steps are completed by considering all non-processed $K_{i}$- or $\mathbf{K}$-values, and Step P3 is considered. By Proposition 6, we obtain a repair $R_{2}$ that satisfies item 2 in the proposition. The proof is therefore complete. $\Box$
@@ 1. Regarding $R_{1}$, we build $\varphi$ as follows: For every $\sigma$ in $\Sigma(\mathcal{AQ})$:
$D_{1}$ $K_{1}$ $A_{1}$ $k_{1}$ $10$ $k^{\prime}_{1}$ $-5$ $D_{2}$ $K_{2}$ $A_{2}$ $k_{2}$ $20$ $k^{\prime}_{2}$ $-1$ $k^{\prime}_{2}$ $30$ $F$ $K_{1}$ $K_{2}$ $M_{1}$ $k_{1}$ $k_{2}$ $-10$ $k_{1}$ $k^{\prime}_{2}$ $2$ $k^{\prime}1$ $k^{\prime}_{2}$ $-100$
Thus $m\_Chase(T)$ contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
[…]
and $T$ has two repairs $R_{1}$ and $R_{2}$. Moreover, the tuples in ${\sf True}(\mathcal{R}_{i})$ ($i=1,2$) whose schemas contain $\mathbf{K}=K_{1}K_{2}$, respectively denoted $R^{\mathbf{K}}_{1}$ and $R^{\mathbf{K}}_{2}$, are:
$R^{\mathbf{K}}_{1}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,-1,2),(k^{\prime}_{1},k^{\prime}_{2},-5,-1,-100)\}$
$R^{\mathbf{K}}_{2}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,30,2),(k^{\prime}_{1},k^{\prime}_{2},-5,30,-100)\}$.
@@
In the following two sections 6.2 and 6.3, we successively introduce our algorithms for computing consistent answers to analytic queries, depending on whether they involve or not a group-by clause. In each case, the main algorithm relies on a procedure, called Compute_aggregate and shown in Algorithm 2, whose role is to scan (either entirely or partially) the m-table $m\_Chase(T)$ and return values that will appear in the answers.
6.2 Analytic Queries with no Group-by Clause
If the query involves no group-by clause, as shown in Algorithm 3, the procedure Compute_aggregate of Algorithm 2 is called to scan the whole m-table $m\_Chase(T)$. When running this call, line 1 in Algorithm 3, the main loop lines 6-32 of Algorithm 2 scans $m\_Chase(T)$ and computes values appearing in the answers. The following proposition shows that Algorithm 3 is correct, except when $aggr=sum$. In other words, if $aggr\neq sum$ Algorithm 3 returns ${\sf C\_ans}(\mathcal{AQ})$. Moreover, it is also shown that ${\sf C\_ans}^{*}(\mathcal{AQ})$ is correctly computed by Algorithm 3 in any case.
@@ 6.2 Analytic Queries with no Group-by Clause
$proc\_change$ // a boolean variable
$proc\_min$ // meant to be equal to the lower bound of ${\sf C\_ans}(\mathcal{AQ})$ if $proc\_change$ has value true
$proc\_max$ // meant to be equal to the upper bound of ${\sf C\_ans}(\mathcal{AQ})$ if $proc\_change$ has value true
$proc\_ans^{*}$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$ if $proc\_change$ has value true
@@
[…]
@@
4:      $proc\_min:=0$ ; $proc\_max:=0$ ; $proc\_ans^{*}:=0$
@@
7:      if $\mathbf{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ then
@@
10:                    $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
11:                    $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
12:               else
13:                    $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i})))$
14:                    $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i})))$
15:           if ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$ then
16:               $proc\_change:={\tt true}$
17:               $proc\_min:=aggr(min_{\sigma},proc\_min)$
18:               $proc\_max:=aggr(max_{\sigma},proc\_max)$
@@
27:                    $proc\_max:=proc\_max+1$

### v2

@@ 5.3 Analytic Queries and their Consistent Answers
In the literature Ullman , an analytic query is a query involving an aggregate operator among countcount, minmin, maxmax, sumsum, and possibly a group-by clause. Formally, given a data warehouse, an analytic query $\mathcal{AQ}$ has one of the following two forms:
@@
where $\varphi$ is the join of the fact table $F$ with all dimension tables $D_{1},\ldots,D_{n}$, $X$ is a relation schema, and for $j=1,\ldots,p,$ aggraggr is an aggregate operator such as countcount, minmin, maxmax, sumsum and $M_{i}$ is a measure attribute.
@@
  • If $\mathcal{AQ}$ involves no group-by clause, ${\sf Ans}(\mathcal{AQ})$ is the value of the aggregation evaluated over the set of all tuples in $\varphi$ satisfying the condition $\Gamma$. We notice that the aggregate may involve no attribute, when expressed as count()count(*).
@@
Now, if the underlying data warehouse does not satisfy all functional dependencies, this traditional semantics of answer to analytic queries has to be revisited. When no dependency of the form $\mathbb{K}\to M_{i}$ is considered, this has been done in ArenasBCHRS03 , in the case of queries with no group-by clause and in FuxmanFM05 when a group-by clause occurs in the query. As shown in DixitK22 , in either case, the ‘repair semantics’ of the consistent answer to an analytic query consists intuitively in producing an interval in which aggregate values fall, when answering the query in any repair. The reader is referred to the forthcoming Section 7 for a more precise relationship between these approaches and our work.
In our approach, we follow the same line regarding analytic queries and their consistent answers, and we emphasize that, as in the previous section, all selection conditions are assumed to be independent.
@@
  • If there exists $R$ in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no tuple over $sch(Q)$ satisfying $\Gamma$, then ${\sf C\_ans}(\mathcal{AQ})={\tt NULL}$ if $aggr\neq count$, and ${\sf C\_ans}(\mathcal{AQ})=0$ if aggraggr is countcount.
@@
  • for every $R$ in ${\sf Rep}(T)$ there exists $d$ in $[glb,lub]$ such that ${\sf Ans}(\mathcal{AQ}^{[R]})$ contains (xd)(x\,d),
  • there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})$ and ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})$ respectively contain $(x,glb)$ and $(x,lub)$. $\Box$
@@
  • In case (A), a NULL answer means that the consistent answer to the query $Q:~{\tt select}$ $\mathbb{K}$ from $T~{\tt where}~\Gamma$ is empty. In this case the aggregate value cannot be computed if the aggregate operator is different than countcount, and if the aggregate operator is countcount, then the expected answer is 0. This explains why we return NULL in the former case.
@@
An alternative way to deal with consistent answers to analytic queries consists of computing first the consistent answer to an appropriate non analytic query and then, of applying the aggregation operators to this answer. More formally, a given analytic query $\mathcal{AQ}:$~select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] is associated with the non analytic query defined by
[…]
and this alternative consistent answer to $\mathcal{AQ}$ is defined as follows.
Definition 6
Let $T$ be a star-table over universe $U$ and let $\mathcal{AQ}$ be an analytic query defined as follows: $\mathcal{AQ}:$~select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$]. The strongly consistent answer to $\mathcal{AQ}$, denoted by ${\sf C\_ans}^{*}(\mathcal{AQ})$, is the answer to the analytic query:
[…]
where ${\sf C\_ans}(Q_{\mathcal{AQ}})$ is the consistent answer to the (non-analytic) query $Q_{\mathcal{AQ}}$ as defined above. $\Box$
We emphasize that, since the select clause of $Q_{\mathcal{AQ}}$ contains $\mathbb{K}$, ${\sf C\_ans}(Q_{\mathcal{AQ}})$ can be effectively computed based on Corollary 1, assuming that m_Chase(T)m\_Chase(T) is available. Moreover, the semantics of such answers are radically different than those defined in Definition 5, simply because strongly consistent answers do not involve intervals, whereas consistent answers do. Moreover, as shown in the following example, aggregate values in the strongly consistent answer may not belong to the corresponding interval of the consistent answer.
@@
In the context of Example 3, $\mathcal{AQ}_{1}:{\tt select}$ $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$) is an example of analytic query with no group-by clause and involving the aggregate sumsum. The expected result is the interval containing all possible sums of $M_{1}$-values among all tuples defined over $K_{1}K_{2}$ and associated with $A_{1}^{1}$-value $a_{1}$ in all repairs. Referring to Figure 4, we have:
@@
Assuming that $m_{1}$, $m^{\prime}_{1}$ and $m^{\prime\prime}_{1}$ are positive numbers and that $m^{\prime}_{1}\leq m^{\prime\prime}_{1}$, we obtain that ${\sf C\_ans}(\mathcal{AQ}_{1})=[glb,lub]$ where $glb=m_{1}$ and $lub=2\cdot m_{1}+m^{\prime\prime}_{1}$.
On the other hand, we have $Q_{\mathcal{AQ}_{1}}:$~select $\mathbb{K},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$), and it can be seen from Figure 4, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{1}})=\{k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$. Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{1})=m_{1}$. Notice that in this case, the value in ${\sf C\_ans}^{*}(\mathcal{AQ}_{1})$ belongs to the interval in ${\sf C\_ans}(\mathcal{AQ}_{1})$.
@@
On the other hand, we have $Q_{\mathcal{AQ}^{\prime}_{1}}:$~select $\mathbb{K},M_{1}$ from $T$ where ($A_{1}^{1}=a^{\prime}_{1}$), and it can be seen from Figure 4, that ${\sf C\_ans}(Q_{\mathcal{AQ}^{\prime}_{1}})=\emptyset$. Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}^{\prime}_{1})={\tt NULL}$.
@@
On the other hand, we have $Q_{\mathcal{AQ}_{2}}:$~select $\mathbb{K},A_{1}^{2},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$). It can be seen from Figure 4, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{2}})=\{k_{1}k_{2}b_{1}m_{1}\}$. Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{2})=\{(b_{1},m_{1})\}$.
It is important to notice that this case illustrates that if $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$ and xmx\,m is in ${\sf C\_ans}^{*}(\mathcal{AQ})$, then it does not always hold that $glb\leq m\leq lub$. Indeed, for $m_{1}=1$, $m^{\prime}_{1}=2$ and $m^{\prime\prime}_{1}=3$, we have $glb=3$, $lub=4$ showing that for $x=b_{1}$, $m\not\in[glb,lub]$. $\Box$
@@ 6 Computing Consistent Answers to Analytic Queries
In this section, we give algorithms for computing consistent answers to analytic queries, first when no group-by clause is present, then when a group-by clause is present. We also consider in this section the case of analytic queries involving a group-by-having clause, whereas the introduction of the clause distinct is the subject of the last sub-section. We emphasize again that the selection conditions in queries are assumed to be independent.
@@ 6.1 Basic Result
[…]
$\bullet$ $\Sigma(\mathcal{AQ})=\{\sigma\in m\_Chase(T)~|~(\mathbb{K}\cup M_{i}\cup sch(\Gamma)\subseteq sch(\sigma))\wedge(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset)\}$
@@
[…]
@@
    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{1}))(t.sch(\Gamma)\not\in Sat(\Gamma))$,
@@
    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{2}))(t.sch(\Gamma)\in Sat(\Gamma))$.
Proof. See Appendix F. $\Box$
@@
$D_{1}$ $K_{1}$ $A_{1}$ $k_{1}$ $10$ $k^{\prime}_{1}$ 5-5 $D_{2}$ $K_{2}$ $A_{2}$ $k_{2}$ 2020 $k^{\prime}_{2}$ 1-1 $k^{\prime}_{2}$ $30$ $F$ $K_{1}$ $K_{2}$ $M_{1}$ $k_{1}$ $k_{2}$ 10-10 $k_{1}$ $k^{\prime}_{2}$ 22 $k^{\prime}1$ $k^{\prime}_{2}$ 100-100
Thus m_Chase(T)m\_Chase(T) contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
$\sigma_{1}=(k_{1})(k_{2})(10)(20)(-10)$, $\sigma_{2}=(k_{1})(k^{\prime}_{2})(10)(-1\,30)(2)$, $\sigma_{3}=(k^{\prime}_{1})(k^{\prime}_{2})(-5)(-1\,30)(-100)$
and $T$ has two repairs $R_{1}$ and $R_{2}$. Moreover, the tuples in ${\sf True}(\mathcal{R}_{i})$ ($i=1,2$) whose schemas contain $\mathbb{K}=K_{1}K_{2}$, respectively denoted $R^{\mathbb{K}}_{1}$ and $R^{\mathbb{K}}_{2}$, are:
$R^{\mathbb{K}}_{1}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,-1,2),(k^{\prime}_{1},k^{\prime}_{2},-5,-1,-100)\}$
$R^{\mathbb{K}}_{2}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,30,2),(k^{\prime}_{1},k^{\prime}_{2},-5,30,-100)\}$.
@@
proc_changeproc\_change // a boolean variable
proc_minproc\_min // meant to be equal to the lower bound of ${\sf C\_ans}(\mathcal{AQ})$ if proc_changeproc\_change has value true
proc_maxproc\_max // meant to be equal to the upper bound of ${\sf C\_ans}(\mathcal{AQ})$ if proc_changeproc\_change has value true
$proc\_ans^{*}$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$ if proc_changeproc\_change has value true
@@
2:    proc_min:=proc\_min:=-dummy ; proc_max:=proc\_max:=+dummy ; $proc\_ans^{*}:=$-dummy// -dummy and +dummy are values not in $adom(M_{i})$ such that for every $m$ in $adom(M_{i})$,// $aggr($-dummy,m)=aggr(,m)=aggr(+dummy,m)=m,m)=m and $\min($-$m)=\max(+dummy$,m)=m,m)=m
@@
4:    proc_min:=0proc\_min:=0 ; proc_max:=0proc\_max:=0 ; $proc\_ans^{*}:=0$
@@
7:    if $\mathbb{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ and $|{\sf tuples}(\sigma(X))|=1$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ then
@@
10:        $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$ ; $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
11:       else
12:        $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i})))$ ; $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i})))$
13:     if ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$ then
14:       $proc\_change:={\tt true}$
15:       if $aggr=count$ then
16:        proc_min:=proc_min+1proc\_min:=proc\_min+1
17:       else
18:        $proc\_min:=aggr(min_{\sigma},proc\_min)$ ; $proc\_max:=aggr(max_{\sigma},proc\_max)$
@@
27:        proc_max:=proc_max+1proc\_max:=proc\_max+1